I need to solve the system of linear equations
x/4 - y/11 = -9/44
x/6 - y/13 = -1/26

Answer by stanbon(75887)

I need to solve the system of linear equations
x/4 - y/11 = -9/44
x/6 - y/13 = -1/26
-------------------------
Multiply thru the 1st equation by 44;
Multiply thru the 2nd equation by 78.
---------------------
11x - 4y = -9
13x - 6y = -3
-----------------
Solve by any method to get:
x = -3
y = -6
=================
Cheers,
Stan H.